# version 0.9

### v5

@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
@@
Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under CC-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust. The corresponding cost is
@@ 2.1.5 Epidemiological Models
Figure 2.5: Trajectory of a continuous-time SIR model.
@@ 2.2.1 Equilibrium of Dynamical Systems
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector. Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
@@ 2.2.3 Quantitative Definitions of Stabilizability
Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, i.e.
@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise. Consider the following dynamical system over the complex numbers
@@
$J_{c}=\{z\in C\ :\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$. In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (which is not $\infty$). The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, i.e. there is no Turing machine that decides it in finite time! (see e.g. [BY08]) The complexity that arises near the boundary of the Julia set is depicted in figure 2.7.
Another compelling visual evidence of the complexity of equilibrium in dynamical systems is the Lorenz attractor. This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
@@
The 3D trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly as depicted in figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This shows that even studying a solution concept itself is a formidable task.
@@
Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under CC-BY-2.5.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e., for every possible starting state, does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
@@
Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-3 polynomial as
@@
and the degree-6 polynomial for the entire formula as
@@ 2.4 Bibliographic Remarks
[Ada20] discusses further aspects of simulations used in the early days of the COVID-19 pandemic. See [SZG+21] for examples of control algorithms used for non-invasive mechanical ventilation.
@@ 2.4. Consider the double integrator system:
The goal is to ensure that the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
@@
[…]
    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of $A-BK$ strictly within the unit circle. Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom left square, and its goal is to reach the $+1$ reward. The blocks $+1,-1$ represent terminal states, which means that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
StartX+1-1
@@
For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
@@
On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
@@
We then describe a formal model to capture this and any other reinforcement learning problem.
@@ 3.2 Markov Decision Processes
To ease the exposition, we start with a review of Markov Chains, a.k.a. Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix that describes the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
@@
Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chance of weather for the next day.
@@
    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ (P^{t})_{ii}>0\right\},$
    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state,
    $\hat{R}(s)=\mathop{\mbox{\bf E}}[R_{t}|S_{t}=s].$
@@
    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards.
    $G_{t}=\sum_{i=0}^{\infty}R_{t+i}\gamma^{i}.$
@@
The value function tells how valuable a state is as measured by the sum of its future discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define
    $R_{sa}=\sum_{s^{\prime}\in S}P_{s,s^{\prime},a}R_{s,s^{\prime},a}.$
@@
A fundamental theorem of reinforcement learning is as follows.
@@ 3.3 The Bellman Equation
    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e.
    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
@@ 3.6 Exercises
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under CC-BY-2.5.
@@ 4.2 Stabilizability of Linear Systems
Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
@@
Suppose that there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
@@ 4.2.1 Efficient Verification of Stabilizability
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions. The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
@@
There exists a matrix ${K}$ such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
@@
Furthermore, given such matrices ${P}$ and ${L}$ satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
@@
Define a block matrix ${M}$ as
@@
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$. Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
@@
Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing ${L}={KP}$.
@@ 4.3 Controllability of Linear Dynamical Systems
Since linear transformations retain containment, we can multiply by ${A}$ and get
@@
Adding ${B}$ to both matrices does not affect the subset relationship.
@@ 4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear time-invariant system given by the transformations $({A,B})$. We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
@@
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@ 5.3.1 Solving the DARE
In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following SDP.

### v6

@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The two dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft:
@@
Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under $C$-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate — each of which is a linear transformation of the state — and applied thrust. The corresponding cost is
@@ 2.1.5 Epidemiological Models
Figure 2.5: Trajectory of a continuous-time SIR model with parameters $\alpha=0,\beta=0.3,\gamma=0.1$.
@@ 2.2.1 Equilibrium of Dynamical Systems
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all A{A}. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector. Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
@@ 2.2.3 Quantitative Definitions of Stabilizability
Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, that is,
@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise. Consider the following dynamical system over the complex numbers:
@@
$J_{c}=\{z\in C\ \colon\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$. In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (that is not $\infty$). The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, that is, there is no Turing machine that can decide it in finite time! (See e.g. [BY08].) The complexity that arises near the boundary of the Julia set is depicted in Figure 2.7.
Another compelling visual illustration of the complexity of equilibrium in dynamical systems is the Lorenz attractor. This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
@@
The three dimensional (3D) trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly, as depicted in Figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This shows that even studying a solution concept itself is a formidable task.
@@
Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under $C$-BY-2.5.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following problem: Given a dynamical system, can we determine if it stabilizable, that is, for every possible starting state, does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
@@
Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-three polynomial as
@@
and the degree-six polynomial for the entire formula as
@@ 2.4 Bibliographic Remarks
[Ada20] discusses further aspects of simulations used in the early days of the COVID-19 pandemic. See [SZG+21] for examples of control algorithms used for noninvasive mechanical ventilation.
@@ 2.4. Consider the double integrator system:
The goal is to ensure that the system reaches the desired state $x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
@@
    $u_{t}=-K(\mathbf{x}_{t}-[x^{*}\;0]^{\top}),$
    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of ABKA-BK strictly within the unit circle. Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of ABKA-BK are real and lie strictly between 1-1 and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom left square, and its goal is to reach the +1+1 reward. The blocks +1,1+1,-1 represent terminal states, which means that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
StartX+1+11-1
@@
For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent +1+1 reward, as given in figure 3.3.
@@
Conversely, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
@@
We next describe a formal model to capture this and any other reinforcement learning problem.
@@ 3.2 Markov Decision Processes
To ease the exposition, we start with a review of Markov Chains, also known as Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix that describes the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as follows:
@@
Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chances of weather for the next day.
@@
    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where d(s_i) = gcd{t ∈N_+ | (P^t)_ii > 0 } , is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state, ^R(s) = E[ R_t | S_t = s] .
@@
    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards. G_t = ∑_i=0^∞ R_t+i γ^i .
@@
The value function says how valuable a state is as measured by the sum of its future discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define R_sa = _s’ S P_s,s’,a R_s,s’,a .
@@
A fundamental theorem of reinforcement learning, which is constructively proven in Section 3.4, is as follows.
@@ 3.3 The Bellman Equation
    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e. P^π_ss’ = ∑_a ∈A P^a_ss’ ⋅π(a|s).
@@ 3.6 Exercises
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under $C$-BY-2.5.
@@ 4.2 Stabilizability of Linear Systems
Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix A{A}, that is,
@@
Suppose that there exists a K{K} such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
@@ 4.2.1 Efficient Verification of Stabilizability
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $I+A^{\top}PA, PP$ satisfies the required conclusions. The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
@@
There exists a matrix K{K} such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
@@
Furthermore, given such matrices P{P} and L{L} satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
@@
Define a block matrix M{M} as
@@
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$. Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices K{K} and P{P} such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices K{K} and P{P} such that
@@
Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing L=KP{L}={KP}.
@@ 4.3 Controllability of Linear Dynamical Systems
Since linear transformations retain containment, we can multiply by A{A} and get
@@
Adding B{B} to both matrices does not affect the subset relationship.
@@ 4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear time-invariant system given by the transformations (A,B)({A,B}). We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
@@
Consider a linear controller K{K} for the LTI system (A,B)({A,B}). If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix A+BK{A+BK} admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@ 5.3.1 Solving the DARE
In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$} to form a quadratic equation in S{S}. Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following semidefinite program.